# DT-DDNN: A Physical Layer Security Attack Detector in 5G RF Domain for CAVs

### v2

@@ Abstract
The Synchronization Signal Block (SSB) is a fundamental component of the 5G New Radio (NR) air interface, crucial for the initial access procedure of Connected and Automated Vehicles (CAVs), and serves several key purposes in the network’s operation. However, due to the predictable nature of SSB transmission, including the Primary and Secondary Synchronization Signals (PSS and SSS), jamming attacks are critical threats. These attacks, which can be executed without requiring high power or complex equipment, pose substantial risks to the 5G network, particularly as a result of the unencrypted transmission of control signals. Leveraging RF domain knowledge, this work presents a novel deep learning-based technique for detecting jammers in CAV networks. Unlike the existing jamming detection algorithms that mostly rely on network parameters, we introduce a double-threshold deep learning jamming detector by focusing on the SSB. The detection method is focused on RF domain features and improves the robustness of the network without requiring integration with the pre-existing network infrastructure. By integrating a preprocessing block that extracts PSS correlation and energy per null resource elements (EPNRE) characteristics, our method distinguishes between normal and jammed received signals with high precision. Additionally, by incorporating of Discrete Wavelet Transform (DWT), the efficacy of training and detection are optimized. A double-threshold double Deep Neural Network (DT-DDNN) is also introduced to the architecture complemented by a deep cascade learning model to increase the sensitivity of the model to variations of signal-to-jamming noise ratio (SJNR). Results show that the proposed method achieves 96.4% detection rate in extra low jamming power, i.e., SJNR between 15 to 30 dB which outperforms the single threshold DNN design with 86.0% detection rate and unprocessed IQ sample DNN design with 83.2% detection rate. Ultimately, the performance of DT-DDNN is validated through the analysis of real 5G signals obtained from a practical testbed, demonstrating a strong alignment with the simulation results.
@@
The evolution in transformative technologies such as Connected and Automated Vehicles (CAVs), the Internet of Things (IoT), and Edge Computing, necessitates the development of next-generation wireless networks to guarantee the Quality of Service (QoS) for communication processes [1]. Given the impact of security attacks on QoS, various defense strategies are proposed to provide detection mechanisms and countermeasures to reduce the impact of these attacks [2][3][4][5]. Among these, machine learning-based defense strategies including decentralized and centralized are rising due to their performance in increasing the robustness of the wireless networks [6]. Moreover, proper operation of critical elements of the network, such as synchronization signal block (SSB), is required to ensure the integrity and security of the communication. SSB is an essential component in 5G wireless communication systems, as it enables device synchronization with the network and access to the services that are available [7]. The SSB functions as a reference signal that provides the establishment and maintenance of connections between mobile devices and the base station (gNB). Assisting user equipment (UE), including smartphones and other devices, in synchronizing with the 5G network is the main function of the SSB. Effective communication requires synchronization, which guarantees that the UE and the base station operate at the same time and frequency. It is crucial that 5G SSBs are robust and resilient to a variety of channel conditions in order to guarantee synchronization that is reliable despite the complexity of the environment. In addition, 5G enables the configuration of SSBs to be more adaptable to various deployment scenarios and facilitates more effective network management.
@@
SSB is transmitted based on predetermined frequency and timing resources in 5G networks [8]. Due to this predictability, an attacker can target these established locations. By identifying the subcarrier spacing and getting synchronized with the cell in the time domain, an attacker is able to extract the cell identity and as a result, locates and targets PSS or SSS which prevents UE from receiving critical signals required for synchronization [9], [10]. By detecting PSS and SSS, and learning about the PCI, the adversary can execute a jamming attack on PBCH [11], [12], [10]. Jamming attacks targeting PBCH in SSB, unable UEs to access necessary information and new connections to cells. Authors in [13] suggested using localization-based detection techniques as a solution, however, these techniques are not efficient in a mobile jammer scenario. Another reason behind the vulnerability of SSB is the lack of encryption through initial access. Unlike user data which is encrypted to ensure confidentiality in 5G networks, control signals such as PSS and SSS are transmitted without encryption [12]. Thus, SSB can potentially be jammed by anyone without the requirement of deciphering or authentication. Furthermore, an attacker which performs jamming attack on SSB does not require to consume high energy to successfully disrupt the communication [14]. Utilizing less jamming power enables the jammer to minimize its impact on the monitoring and detection systems as most of the detection systems rely on identifying anomalous high-energy patterns. On the other hand, it helps the attacker to perform the attack with simple and inexpensive equipment. Traditional jamming detection methods rely on the received signal intensity, or the performance of the network [15]. The signal level detection methods are incapable of maintaining the detection accuracy in high signal-to-jamming and noise ratio (SJNR) scenarios, and are mostly efficient in constant jamming cases [16], [13]. The detection techniques based on the network metrics such as packet delivery ratio (PDR) or bit error rate (BER) may be unsuccessful when facing advanced jammers utilizing selective or intelligent techniques which may not have a significant impact on the error rate or performance metrics.
@@
This research presents a jamming detection technique for 5G networks based on a deep learning approach operating solely in the RF domain which eliminates the need for the jamming detector to be an entity of the network. A key discriminator that sets our work apart from existing jamming detection algorithms is its focus on the analysis of 5G SSB using the design of a double threshold double DNN model (DT-DDNN) to enable the system to detect both smart and barrage jammers effectively. This expands the algorithm’s applicability and impact by covering a wider range of jamming scenarios. To enhance the precision of the detection algorithm and reduce its time and energy consumption, a preprocessing block is integrated into the design in which correlation and EPNRE characteristics of the PSS are extracted. This step provides a more precise discrimination between normal and jammed receive signals. As the deep learning algorithms require advanced computational resources, we incorporate the Discrete Wavelet Transform (DWT) into our design. The DWT magnifies the features that were extracted during the preprocessing phase, and provides a more detailed representation of the signal attributes. This not only improves the performance of the detection model but also optimizes the training procedure. A significant enhancement to our architecture is the integration of double threshold DNN which has been used to improve precision in the scenarios indicated by higher SJNR values in which the detector faces challenges to make a final decision. The additional DNN is supported by a deep cascade learning model to increase the sensitivity of the design in a high SJNR regime. Therefore, the main contributions of this study are summarized as follows.
@@
Utilize over-the-air 5G signal features without the need for higher layer KPIs such as Block Error Rate (BLER), Bad Packet Ratio (BPR), throughput, and other metrics in higher layers of 5G protocol stack. This enables the jamming detection module to be implemented independently in 5G networks,.i.e., the proposed method does not need to be implemented on 5G network entities such as gNB, UE, or 5GC.
@@ II Related Work
The primary aim of physical layer security (PLS) is to fulfill the security demands of wireless networks by leveraging the unique characteristics of the physical layer [17]. There has been significant research dedicated to developing techniques for detecting attacks that specifically target the physical layer of wireless communication systems. In the context of the vulnerabilities associated with 5G Non-Orthogonal Multiple Access (NOMA), [18] discusses challenges that arise when a near-end user (NU) engages in eavesdropping activities. To address this, modifications to the serial interference cancellation (SIC) process through cooperative jamming have been suggested. These modifications aim to devise optimal power allocation strategies tailored to various situations, enhancing the secrecy rates without compromising the data throughput of legitimate users, ensuring it remains within acceptable limits.
[18] discusses the challenges of 5G Non-Orthogonal Multiple Access (NOMA) while a near-end user (NU) performs an eavesdropping attack. The authors proposed modifying the order of serial interference cancellation (SIC) using cooperative jamming. Optimal power allocation strategies are designed for different scenarios to improve secrecy rates while maintaining the data rate of authorized users within the specific threshold. Authors in [19] analyze how the reusability of resources compromises the physical layer security in multiple-input multiple-output NOMA (MIMO-NOMA) scenario in cognitive radio network (CRN). Thus, a transmit-zero-forcing beamforming (ZFBF) method is proposed alongside signal alignment in an environment while primary and secondary users share cells. Furthermore, the eigen beamforming method is used to maximize the signal-to-information-leakage-plus-noise ratio (SLNR) and therefore improve the physical layer security. AN innovative minimal angle-difference user pairing scheme is provided in [20] to mitigate the impact of eavesdroppers in the millimeter wave (mmWave) NOMA network. This work also presents and assesses two secrecy beamforming models, NMRT and CMRT, which increase the secrecy performance by taking advantage of the spatial correlation between the user and eavesdropper. The efficacy of physical layer security of the integration of reconfigurable intelligent surface (RIS) and Radio Frequency-Underwater Optical Wireless Communication (UOWC) in a scenario including an eavesdropper is investigated in [21]. [22] increases the secrecy rate of mmWave vehicular communication in response to eavesdropping attacks by using a multiantenna transmission technique and a robust beamforming model. The beamforming vectors are optimized to guarantee the QoS for authorized users and maximize the secrecy rate at the same time. Authors in [23] present a novel anti-jamming technique based on federated deep reinforcement learning (DRL) which is based on a joint beamforming and power allocation optimization problem. By using the Markov decision process (MDP), the optimization problem is transformed into a multi-agent reinforcement learning (MARL) problem. The main purpose of the anti-jamming method is to improve the achievable rate of femtocell users (FUs) during a jamming attack targeting 5G Heterogeneous Networks (Het-Nets). [24] focuses on attack identification in air-to-ground communication links using deep attention recognition (DAtR) which detects security attacks using a small deep network embedded in legitimate unmanned aerial vehicles (UAVs). The system employs two metrics of received signal strength indicator (RSSI) and signal to interference plus noise ratio (SINR) to identify attacks in different scenarios such as non-line-of-sight (NLoS), line-of-sight (LoS), or a combination of both. In [25] a jamming detection and defense strategy is proposed by implementing pseudo-random blanking of resource elements within OFDM symbols based on statistical hypothesis testing. The work is focused on smart jammers which maximize their impacts by reducing spectral efficiency and BLER values, and minimize their detection probability. Through the detection algorithm, certain subcarriers are left blank across OFDM symbols in a pseudo-random fashion. Adopting a pseudo-random algorithm that determines the blanking pattern makes the system unpredictable and increases its robustness to jamming. The suggested approach incorporates the downlink data transmission system of 5G without necessitating adjustments to the current infrastructure. [16] and [26] present a novel approach that employs the sensitivity of Error Vector Magnitude (EVM) to detect the presence of tone and chirp jammers when the EVM value reaches a predetermined threshold. Besides the high sensitivity and minimal complexity, this approach offers spectral information about the jammer and affected frequency bands. [27] describes an adaptable, multidimensional strategy for detecting and classifying jamming attacks, considering power levels and frequency band variations. The detection technique is based on supervised learning and receives metrics such as channel quality indicator (CQI), bit rates, packet rates, power headroom, and power headroom. The dataset is generated using a 5G NR non-standalone (NSA) testbed from cross-layer key performance indicator (KPI) data. By combining instantaneous discriminative models along with sequential time series-based models, adaptability is obtained. The use of Reservoir Computing (RC) helps to achieve high accuracy and low false alarm rates while following the latency limits of real-time scenarios and requiring minimal computational and memory resources. The primary focus of [28] is the utilization of ensemble learning and the XGBOOST-ensemble learning combination as a machine learning-based jamming detection in C-RAN. Authors use the WSN-DS database to assess the performance of various machine learning algorithms. The WSN-DS database includes 374,661 samples as chosen features that encapsulate the behavioral patterns of network traffic under normal and jammed conditions. Data preprocessing is performed to set up the dataset for analysis by separating the independent and dependent variables. The feature extraction process is carried out using the XGBOOST algorithm which emphasizes important patterns within the data. Model training is then developed based on cross-validation with various algorithms to identify the most effective model for detection. The model is finally tested and validated using an independent dataset. [15] emphasizes the need for real-time detection and mitigation techniques. Hence, the authors make a contribution by introducing a technique for real-time detection of jamming attacks, utilizing the Hoeffding decision tree. This machine learning methodology facilitates real-time processing and addresses limitations encountered in conventional decision tree models. The jamming node in the simulation uses numerous antennas to simulate a smart jammer and a comprehensive dataset under congestion and normal conditions is created to feed to the Hoeffding algorithm. [29] presents a multi-stage machine learning-based intrusion detection system (ML-IDS) specifically designed for 5G Cloud Radio Access Network (C-RAN) which is designed to detect constant, random, deceptive, and reactive jammers. A Multilayer Perceptron (MLP), a deep learning algorithm, and a Kernelized Support Vector Machine (KSCM) compose the proposed ML-IDS. The MLP is implemented as the main classifier, and in situations where it produces a false negative, the KSVM functions as a supplementary verification mechanism to enhance the precision of jamming attack detection. The ML-IDS is incorporated into the Base-Band Unit (BBU) pool, an essential element of the C-RAN responsible for monitoring spectrum distribution and user information. This enables the ML-IDS to efficiently manage traffic between base stations. The integration of supervised and unsupervised learning methods for jamming detection is introduced in [30]. Known jamming attacks are detected using a supervised learning model while an unsupervised anomaly detection method using auto-encoders is applied for unknown jamming types. To enhance the accuracy of the detection algorithm, address the computational complexity, and causality analysis of KPI degradation, the Bayesian Network Model (BNM) is integrated into the design. The learning-based model is executed based on network parameters such as bitrate, packet rate, retransmission rate, and CQI. [12] concentrates on intelligent PBCH jamming (PBCH-IJ) attack in 5G NR which disrupts the MIB decoding by applying a sniffing attack to extract PCI information. Based on the observations, authors concluded that the PBCH-IJ has a substantial effect on the principal direction in low mobility scenarios. Thus, the introduced counteract is based on the Principal Direction of PBCH Dominant Space (PDPDS) and is executed at user side. To identify anomalies that can detect the jamming attack, sparse principal component analysis (SPCA) is used in the design along with an adaptive detection threshold. The study in [31] implements a machine learning-based detection technique using NS-3 simulator for constant, random, and reactive jammers. Received signal strength, carrier sense time, noise, and PDR are used as jamming detection metrics. The detection method is tested based on three different learning algorithms K nearest neighbor (KNN), decision tree, and random forest with the highest accuracy of 81%. This work distinguishes itself from the aforementioned research by focusing on smart SSB jamming detection in the RF domain using signal features of 5G network. The detection technique is a machine learning-based model using two DNN blocks one with a single threshold and the second one with double threshold characteristics to be able to detect the existence of the jammer in high SJNR conditions. To further enhance the performance of the DNN model deep cascade learning approach is adapted by the second DNN block. A summary of the current literature is provided in Table I.
@@
| Ref. | Technical Approach / Network Model | Feature or Parameter | Attack Type | Real data test |
@@
| [22] | Analytical (optimal beamforming vector)/ mmWave V2V | CSI222channel state information, Transmit Power | Eavesdropper | ✗ |
| [23] | ML(DRL)/5G Het-Nets | channel, beamforming vectors | intermittent, constant, adaptive, reactive | ✗ |
@@
| [16][26] | Statistical (EVM)/5G | EVM | Tone, Chirp Jammer | ✗ |

### v3

@@ Abstract
The Synchronization Signal Block (SSB) is a fundamental component of the 5G New Radio (NR) air interface, crucial for the initial access procedure of Connected and Automated Vehicles (CAVs) and serves several key purposes in the network’s operation. However, due to the predictable nature of SSB transmission, including the Primary and Secondary Synchronization Signals (PSS and SSS), jamming attacks are critical threats. Leveraging radio frequency (RF) domain knowledge, this work presents a novel deep learning-based technique focusing on SSB for detecting jammers in CAV networks without changing pre-existing infrastructure. By integrating a preprocessing block that extracts PSS correlation and energy per null resource elements (EPNRE) characteristics, our method distinguishes between normal and jammed received signals with high precision. Additionally, by incorporating Discrete Wavelet Transform (DWT), the efficacy of training and detection are optimized. A double-threshold double Deep Neural Network (DT-DDNN) is also introduced to the architecture complemented by a deep cascade learning model to increase the sensitivity of the model to variations of signal-to-jamming noise ratio (SJNR). Results show that the proposed method achieves 96.4% detection rate in extra low jamming power, i.e., SJNR between 15 to 30 dB which outperforms the single threshold DNN design with 86.0% detection rate and unprocessed IQ sample DNN design with 83.2% detection rate. Ultimately, the performance of DT-DDNN is validated through the analysis of real 5G signals obtained from a practical testbed, demonstrating a strong alignment with the simulation results.
@@
The evolution in transformative technologies such as Connected and Automated Vehicles (CAVs), the Internet of Things (IoT), and Edge Computing, necessitates the development of next-generation wireless networks to guarantee the Quality of Service (QoS) for communication processes [1]. Given the impact of security attacks on QoS, various defense strategies, particularly machine learning-based techniques, are proposed to enhance network robustness [2][3][4][5][6]. Moreover, proper operation of critical elements of the network, such as synchronization signal block (SSB), is crucial for maintaining device-network synchronization and ensuring service integrity [7]. The SSB functions as a reference signal that guarantees that the user equipment (UE) and the base station operate at the same time and frequency. 5G SSBs must be robust and resilient to a variety of channel conditions to guarantee synchronization that is reliable despite the complexity of the environment. In addition, 5G enables the configuration of SSBs to be more adaptable to various deployment scenarios and facilitates more effective network management.
@@
SSB is transmitted based on predetermined frequency and timing resources in 5G networks [8]. Thus, by identifying the subcarrier spacing and getting synchronized with the cell in the time domain, an attacker can extract the cell identity and as a result, locate and target Primary Synchronization Signal (PSS) or Secondary Synchronization Signal (SSS) which prevents UE from receiving critical signals required for synchronization [9], [10]. Detecting PSS and SSS results in the extraction of Physical Cell Identity (PCI) which can lead to a jamming attack on the Physical Broadcast Channel (PBCH) and prevent UEs from accessing necessary information and new connections to cells [11], [12], [10]. Furthermore, unlike encrypted user data, control signals such as PSS and SSS are transmitted unencrypted [12]. Thus, SSB can potentially be jammed by anyone without the requirement of deciphering or authentication. Jamming attacks on SSB require less jamming power to disrupt the communication [13] enabling the jammer to reduce the detection probability as most of the detection systems rely on identifying anomalous high-energy patterns. Meanwhile, it helps the attacker to perform the attack with simple and inexpensive equipment. Traditional jamming detection methods rely on the received signal intensity, or the performance of the network [14]. The signal level detection methods are incapable of maintaining the detection accuracy in high signal-to-jamming and noise ratio (SJNR) scenarios, and are mostly efficient in constant jamming cases [15], [16]. The detection techniques based on the network metrics such as packet delivery ratio (PDR) or bit error rate (BER) may be unsuccessful when facing advanced jammers utilizing selective or intelligent techniques which may not have a significant impact on the error rate or performance metrics.
@@
To address the security requirements of wireless networks, it is crucial to investigate solutions that operate effectively within the limitations of current infrastructure. This research presents a jamming detection technique for 5G networks based on a deep learning approach operating solely in the radio frequency (RF) domain, employing characteristics of RF signals such as energy patterns and signal correlations without the need for decryption or deep integration into network layers. This approach not only simplifies implementation by avoiding infrastructural modifications but also enables real-time security monitoring required for applications such as CAVs. A key discriminator that sets our work apart from existing jamming detection algorithms is its focus on the analysis of 5G SSB using the design of a double threshold double Deep Neural Network model (DT-DDNN) to enable the system to detect both smart and barrage jammers effectively. This expands the algorithm’s applicability and impact by covering a wider range of jamming scenarios. A critical component of our detection system is the preprocessing block designed to extract features from the received signal that are indicative of jamming activities. This block focuses on the correlation of the PSS and the measurement of Energy per Null Resource Elements (EPNRE), which are significant indicators of potential jamming. The PSS correlation provides a robust method for detecting synchronization disruptions commonly resulting from jamming signals, while the EPNRE metric helps identify unusual energy patterns in the SSB subcarriers where no transmission should occur in a normal transmission scenario. As the deep learning algorithms require advanced computational resources, we incorporate the Discrete Wavelet Transform (DWT) into our design. The DWT magnifies the features that were extracted during the preprocessing phase, and provides a more detailed representation of the signal attributes. This not only improves the performance of the detection model but also optimizes the training procedure. A significant enhancement to our architecture is the integration of double threshold Deep Neural Network (DNN) which has been used to improve precision in the scenarios indicated by higher SJNR values in which the detector faces challenges to make a final decision. The additional DNN is supported by a deep cascade learning model to increase the sensitivity of the design in a high SJNR regime. Therefore, the main contributions of this study are summarized as follows.
@@
Utilize over-the-air 5G signal features without the need for higher layer key performance indicators (KPIs) such as Block Error Rate (BLER), Bad Packet Ratio (BPR), throughput, and other metrics in higher layers of 5G protocol stack. This enables the jamming detection module to be implemented independently in 5G networks,.i.e., the proposed method does not need to be implemented on 5G network entities such as gNodeB (gNB), UE, or 5G core.
@@ II Related Work
| Ref. | Technical Approach / Network Model | Feature or Parameter | Attack Type | Testbed |
@@
| [21] | Analytical (optimal beamforming)/ mmWave V2V | CSI222channel state information, Transmit Power | Eavesdropper | ✗ |
| [22] | ML(DRL)/5G Het-Nets | channel, beamforming vectors | constant, adaptive, reactive | ✗ |
@@
| [15][24] | Statistical (EVM)/5G | EVM | Tone, Chirp Jammer | ✓ |
@@
The primary aim of physical layer security (PLS) is to improve wireless network security by leveraging the unique characteristics of the physical layer [32]. Significant research is focused on identifying attacks on the physical layer of wireless networks. Regarding 5G Non-Orthogonal Multiple Access (NOMA) vulnerabilities, [17] discusses the issue of near-end user (NU) eavesdropping activities. Solutions include modifying serial interference cancellation (SIC) process through cooperative jamming to develop optimal power allocation strategies thereby improving the secrecy rates while maintaining the data throughput of legitimate users within acceptable limits. Authors in [18] analyze how resource reusability compromises physical layer security in multiple-input multiple-output NOMA (MIMO-NOMA) suggesting zero-forcing beamforming (ZFBF) with signal alignment and eigen beamforming method is used to maximize the signal-to-information-leakage-plus-noise ratio (SLNR) and therefore improve the physical layer security. A minimal angle-difference user pairing scheme in millimeter wave (mmWave) NOMA network along with two secrecy beamforming models taking advantage of the spatial correlation between the user and attacker is provided in [19] to increase the secrecy rate in response to an eavesdropper. The efficacy of physical layer security of the integration of reconfigurable intelligent surface (RIS) and Radio Frequency-Underwater Optical Wireless Communication (UOWC) in a scenario including an eavesdropper is investigated in [20]. The research in [21] develops a multiantenna transmission and beamforming model to increases the secrecy rate of mmWave vehicular communication. Authors in [22] present a novel anti-jamming technique based on federated deep reinforcement learning (DRL) which is based on a joint beamforming and power allocation optimization problem. The study in [23] focuses on attack identification in air-to-ground communication links using deep attention recognition (DAtR) which detects security attacks using a small deep network embedded in legitimate unmanned aerial vehicles (UAVs). The system employs two metrics of received signal strength indicator (RSSI) and signal to interference plus noise ratio (SINR) to identify attacks in different scenarios such as non-line-of-sight (NLoS), line-of-sight (LoS), or a combination of both. In [25] a jamming detection and defense strategy is proposed by implementing pseudo-random blanking of resource elements within orthogonal frequency-devision mutiplexing (OFDM) symbols based on statistical hypothesis testing. The work is focused on smart jammers which maximize their impacts by reducing spectral efficiency and BLER values, and minimize their detection probability. Through the detection algorithm, certain subcarriers are left blank across OFDM symbols in a pseudo-random fashion. Adopting a pseudo-random algorithm that determines the blanking pattern makes the system unpredictable and increases its robustness to jamming. The suggested approach incorporates the downlink data transmission system of 5G without necessitating adjustments to the current infrastructure. The authors in [15] and [24] present a novel approach that employs the sensitivity of Error Vector Magnitude (EVM) to detect the presence of tone and chirp jammers when the EVM value reaches a predetermined threshold. Besides the high sensitivity and minimal complexity, this approach offers spectral information about the jammer and affected frequency bands. In [26], an adaptable, multidimensional strategy is introduced for detecting and classifying jamming attacks, considering power levels and frequency band variations. The detection technique is based on supervised learning and receives metrics such as channel quality indicator (CQI), bit rates, packet rates, and power headroom. The primary focus of [27] is the utilization of ensemble learning and the XGBOOST-ensemble learning combination as a machine learning-based jamming detection in C-RAN. Authors use the WSN-DS database to assess the performance of various machine learning algorithms. The WSN-DS database includes 374,661 samples as chosen features that encapsulate the behavioral patterns of network traffic under normal and jammed conditions. Data preprocessing is performed to set up the dataset for analysis by separating the independent and dependent variables. The feature extraction process is carried out using the XGBOOST algorithm which emphasizes important patterns within the data. The study in [14] emphasizes the need for real-time detection and mitigation techniques. Hence, the authors make a contribution by introducing a technique for real-time detection of jamming attacks, utilizing the Hoeffding decision tree. This machine learning methodology facilitates real-time processing and addresses limitations encountered in conventional decision tree models. The work in [28] presents a multi-stage machine learning-based intrusion detection system (ML-IDS) specifically designed for 5G Cloud Radio Access Network (C-RAN) which is designed to detect constant, random, deceptive, and reactive jammers. A Multilayer Perceptron (MLP), a deep learning algorithm, and a Kernelized Support Vector Machine (KSCM) compose the proposed ML-IDS. The integration of supervised and unsupervised learning methods for jamming detection is introduced in [29]. Known jamming attacks are detected using a supervised learning model while an unsupervised anomaly detection method using auto-encoders is applied for unknown jamming types. The learning-based model is executed based on network parameters such as bitrate, packet rate, retransmission rate, and CQI. In [12], the focus is on intelligent PBCH jamming (PBCH-IJ) attack in 5G NR which disrupts the Master Information Block (MIB) decoding by applying a sniffing attack to extract PCI information. To identify anomalies that can detect the jamming attack, sparse principal component analysis (SPCA) is used in the design along with an adaptive detection threshold. The study in [30] implements a machine learning-based detection technique using Network Simulator 3 (NS-3) simulator for constant, random, and reactive jammers. Received signal strength, carrier sense time, noise, and PDR are used as jamming detection metrics. The detection method is tested based on three different learning algorithms K nearest neighbor (KNN), decision tree, and random forest with the highest accuracy of 81%. This work distinguishes itself from the aforementioned research by focusing on smart SSB jamming detection in the RF domain using signal features of a 5G network. In this work, the smart SSB jammer detects the synchronization signal information on the 5G resource grid (including time, frequency, and pattern) and disrupts the corresponding resource blocks, making its detection highly challenging [33]. The detection technique is a machine learning-based model using two DNN blocks one with a single threshold and the second one with double threshold characteristics to be able to detect the existence of the jammer in high SJNR conditions. To further enhance the performance of the DNN model deep cascade learning approach is adapted by the second DNN block. A summary of the current literature is provided in Table I.